b. Two people ? and ? are independently shoot at the same target once, and the hit rate is 0.6 and 0.5, respectively. Given that the target is hit, what is the probability that it is hit by person ??
Answer #1

P(A) = 0.6

P(B) = 0.5

P(target is hit) = P(A) + P(B) - P(A) * P(B)

= 0.6 + 0.5 - 0.6*0.5

= 0.8

P(A|target is hit) = P(A)/P(target is hit) = 0.6/0.8 = 0.75

Required probability = 0.75